OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (-7, 5pi/3)

terenzreignz (terenzreignz):

ahh the easiest of them all :) Given the polar coordinates \((\color{green}r,\color{orange}\theta)\) the rectangular coordinates are: \[\Large (\color{red}x,\color{blue}y )= ( \ (\color{green}r\cos(\color{orange}\theta)\quad , \quad\color{green}r\sin(\color{orange}\theta) \ )\] Simple, no? :P

OpenStudy (anonymous):

Very simple! Just inputted values and came up with -7/2, 7sqrt3/2

OpenStudy (anonymous):

Beast man. Thanks so much!

terenzreignz (terenzreignz):

That was fast... great work :)

OpenStudy (anonymous):

:D